Floor plan of the taylors kitchen.family are looking at options for tiling the floor.kitchen 6m×3m.(rectangle).how many tiles each measuring 100cm×100cm would be needed to cover the entire kichen floor?.2) if the 1m×1m tile come in boxes of five,many boxes would be needed?.
Answer by mananth(16946)

Area of Kitchen = L*W => 6*3=18 m^2
Area of tile = 1 *1 = 1m^2
Number of tiles required 18/1= 18 tiles
18 tiles are required
5 per box
so 18/5 = 3.6 boxes
So 4 boxes will be required